# thanks: These two authors contribute equally in this workthanks: These two authors contribute equally in this work

### v1

@@ II Derivations and Supplemental Results of Superconducting Mean-Field Analysis
The NNN attraction $H_{V_{2}}$ generally leads to a SC phase, while the onsite Hubbard repulsion $H_{U}$ typically favors magnetic ordered phase ($120^{\circ}$ AFM order in this case). Since the SC phase and magnetic ordered phase typically do not coexist, here we assume vanishing magnetic order in the SC mean-field analysis. Thus, according to Eq. (9) in the main text, the mean-field decoupling of $H_{U}$ only contributes an additional sublattice potential term $\frac{1}{2}\sum_{i\alpha}U_{\alpha}n_{i\alpha}$.
@@
Pairing Symmetry Irreducible Representation $f(\delta_{1})$ $f(\delta_{2})$ $f(\delta_{3})$
Mixed $s$-$f$ wave $A_{1}$ $1$ $1$ $1$
Mixed $p_{x}\pm ip_{y}$ $d_{xy}\mp id_{x^{2}-y^{2}}$ wave $E$ $1$ $e^{\pm i\frac{2\pi}{3}}$ $e^{\mp i\frac{2\pi}{3}}$
Mixed $p_{x}$ $d_{xy}$ wave $E$ 00 $\frac{\sqrt{3}}{2}$ $-\frac{\sqrt{3}}{2}$
Mixed $p_{y}$ $d_{x^{2}-y^{2}}$ wave $E$ $1$ $-\frac{1}{2}$ $-\frac{1}{2}$
@@
Pairing Symmetry Irreducible Representation $f(\delta_{1})$ $f(\delta_{2})$ $f(\delta_{3})$
$s$ wave $A_{1}$ $1$ $1$ $1$
Mixed $p_{x}\pm ip_{y}$ $d_{xy}\mp id_{x^{2}-y^{2}}$ wave $E$ $1$ $e^{\pm i\frac{2\pi}{3}}$ $e^{\mp i\frac{2\pi}{3}}$
@@
Taking $U_{A}=U_{B}=35$ meV, $U_{C}=25$ meV, and $V_{2}=10$ meV as in the main text for the NNN Hamiltonian $H$ (see Eq. (S2)), and $V_{1}=10$ meV for the NN Hamiltonian $H^{\prime}$ (see Eq. (S8)), we perform the mean-field calculations for all pairing symmetries classified in Tab. S1 and Tab. S2. The resulting energy gain per hole for NNN and NN Hamiltonians are illustrated in Fig. S2(a) and (b) respectively.
A direct comparison of the energy gains $\Delta E$ reveals that NNN pairings are approximately $2\sim 3$ orders of magnitude stronger than NN pairings for same attraction strength $V_{1}=V_{2}$. We thus conclude that the NNN pairings should be more dominant than the NN pairings, justifying our main focus of the NNN pairings in the main text. This result is actually straightforward to understand. Due to the time reversal symmetry $\mathcal{T}$ in both $H$ and $H^{\prime}$, the Cooper pairs will form between time-reversal-related states. As the time reversal symmetry $\mathcal{T}$ does not alter sublattice index $\alpha$, and given that the Fermi surface have relatively strong sublattice polarization property (as shown in Fig. 2(a) in the main text), the system is expected to favor intra-sublattice pairings (e.g. NNN pairings) over inter-sublattice pairings (e.g. NN pairings). Considering the layer polarization property of the A and B sublattices (the main components of the Fermi surface), it is equivalent to say that the tWSe2 system favors intra-layer pairings.
@@
Returning to the non-nematic cases for NNN pairings, one can see that at small $\mathcal{V}_{z}$ (as discussed in the main text) or large enough $\mathcal{V}_{z}$ regime, the chiral mixed $p_{x}\pm ip_{y}$ $d_{xy}\mp id_{x^{2}-y^{2}}$-wave symmetry is the most dominant, while at intermediate $\mathcal{V}_{z}$, the mixed $s$-$f$-wave symmetry becomes the strongest one. To understand this, we plot the magnitude of the momentum space form factor for NNN pairings with different partial wave components in Fig. S3. As shown in Fig. S3, the $s$-wave component has local maximum at $\gamma$ and $\kappa_{\pm}$ points of Brillouin zone(BZ) while having ring-shaped minimum in between; the $p_{x}\pm ip_{y}$-wave component is large except at $\gamma$ points and the BZ boundary; the $d_{xy}\pm id_{x^{2}-y^{2}}$-wave component has maximum at $m$ points while having minimum at $\gamma$ and $\kappa_{\pm}$ points; and the $f$ wave component has maximum at $\kappa_{\pm}$ points but having minimum nodal line along $\gamma$-$m$ directions.
@@
III Derivations of Magnetic Mean-Field Analysis
@@
Then, we can perform the standard mean-field calculations as follow. We first construct and diagonalize the mean-field Hamiltonian in Eq. (S10), and filled the lowest energy states until $\nu=-1$. We then use Eq. S11 to update the order parameters for next iteration. Such process is repeated until self-consistency is achieved, and the numerical results are detailed in the main text.
@@
$H_{U}\approx-\sum_{k\alpha}U_{\alpha}\left(m_{\alpha}c_{\alpha\vec{k}\downarrow}^{\dagger}c_{\alpha\vec{k}\uparrow}+m_{\alpha}^{*}c_{\alpha\vec{k}\uparrow}^{\dagger}c_{\alpha\vec{k}\downarrow}\right)+N\sum_{\alpha}U_{\alpha}m_{\alpha}^{2}+\frac{1}{2}\sum_{k\alpha\sigma}U_{\alpha}n_{k\alpha\sigma},$ (S12)

### v2

@@ VI Derivations and Supplemental Results of Superconducting Mean-Field Analysis
The NNN attraction $H_{V_{2}}$ generally leads to a SC phase, while the onsite Hubbard repulsion $H_{U}$ typically favors magnetic ordered phase ($120^{\circ}$ AFM order in this case). Since the SC phase and magnetic ordered phase typically do not coexist, here we assume vanishing magnetic order in the SC mean-field analysis. Thus, according to Eq. (III) in the main text, the mean-field decoupling of $H_{U}$ only contributes an additional sublattice potential term $\frac{1}{2}\sum_{i\alpha}U_{\alpha}n_{i\alpha}$.
@@
Pairing Symmetry
Irreducible
Representation
$f(\delta_{1})$ $f(\delta_{2})$ $f(\delta_{3})$
Mixed $s$-$f$ wave $A_{1}$ 11 11 11
Mixed $p_{x}\pm ip_{y}$ - $d_{xy}\mp id_{x^{2}-y^{2}}$ wave $E$ 11 $e^{\pm i\frac{2\pi}{3}}$ $e^{\mp i\frac{2\pi}{3}}$
Mixed $p_{x}$ - $d_{xy}$ wave $E$ 0 $\frac{\sqrt{3}}{2}$ $\frac{\sqrt{3}}{2}$
Mixed $p_{y}$ - $d_{x^{2}-y^{2}}$ wave $E$ 11 $\frac{1}{2}$ $\frac{1}{2}$
@@
Pairing Symmetry
Irreducible
Representation
$f(\delta_{1})$ $f(\delta_{2})$ $f(\delta_{3})$
$s$ wave $A_{1}$ 11 11 11
Mixed $p_{x}\pm ip_{y}$ - $d_{xy}\mp id_{x^{2}-y^{2}}$ wave $E$ 11 $e^{\pm i\frac{2\pi}{3}}$ $e^{\mp i\frac{2\pi}{3}}$
@@
Taking $U_{A}=U_{B}=U_{C}=37.5$ meV, $V_{2}=10$ meV as in the main text for the NNN Hamiltonian $H$ (see Eq. (S14)), and $V_{1}=10$ meV for the NN Hamiltonian $H^{\prime}$ (see Eq. (S20)), we perform the mean-field calculations for all pairing symmetries classified in Tab. S1 and Tab. S2. The resulting energy gain per hole for NNN and NN Hamiltonians are illustrated in Fig. S6(a) and (b) respectively.
A direct comparison of the energy gains $\Delta E$ reveals that NNN pairings are approximately $2\sim 3$ orders of magnitude stronger than NN pairings for same attraction strength $V_{1}=V_{2}$. We thus conclude that the NNN pairings should be more dominant than the NN pairings, justifying our main focus of the NNN pairings in the main text. This result is actually straightforward to understand. Due to the time-reversal symmetry $\mathcal{T}$ in both $H$ and $H^{\prime}$, the Cooper pairs will form between time-reversal-related states. As the time-reversal symmetry $\mathcal{T}$ does not alter sublattice index $\alpha$, and given that the Fermi surface have relatively strong sublattice polarization property (as shown in Fig. 2(a) in the main text), the system is expected to favor intra-sublattice pairings (e.g. NNN pairings) over inter-sublattice pairings (e.g. NN pairings). Considering the layer polarization property of the A and B sublattices (the main components of the Fermi surface), it is equivalent to say that the tWSe2 system favors intra-layer pairings.
@@
Returning to the non-nematic cases for NNN pairings, one can see that at small $\mathcal{V}_{z}$ (as discussed in the main text) or large enough $\mathcal{V}_{z}$ regime (where the energies of different pairing symmetries are close), the chiral mixed $p_{x}\pm ip_{y}$ - $d_{xy}\mp id_{x^{2}-y^{2}}$-wave symmetry is the most dominant, while at intermediate $\mathcal{V}_{z}$, the mixed $s$-$f$-wave symmetry becomes the strongest one. To understand this, we plot the magnitude of the momentum space form factor for NNN pairings with different partial wave components in Fig. S7. As shown in Fig. S7, the $s$-wave component has local maximum at $\gamma$ and $\kappa_{\pm}$ points of Brillouin zone (BZ) while having ring-shaped minimum in between; the $p_{x}\pm ip_{y}$-wave component is large except at $\gamma$ points and the BZ boundary; the $d_{xy}\pm id_{x^{2}-y^{2}}$-wave component has maximum at $m$ points while having minimum at $\gamma$ and $\kappa_{\pm}$ points; and the $f$ wave component has maximum at $\kappa_{\pm}$ points but having minimum nodal line along $\gamma$-$m$ directions.
@@
VII Derivations of Magnetic Mean-Field Analysis
@@
Then, we can perform the standard mean-field calculations as follow. We first construct and diagonalize the mean-field Hamiltonian in Eq. (S22), and filled the lowest energy states until $\nu=-1$. We then use Eq. (S23) to update the order parameters for next iteration. Such process is repeated until self-consistency is achieved, and the numerical results are detailed in the main text.
@@
$H_{U}\approx-\sum_{k\alpha}U_{\alpha}\left(m_{\alpha}c_{\alpha k\downarrow}^{\dagger}c_{\alpha k\uparrow}+m_{\alpha}^{*}c_{\alpha k\uparrow}^{\dagger}c_{\alpha k\downarrow}\right)+N\sum_{\alpha}U_{\alpha}m_{\alpha}^{2}+\frac{1}{2}\sum_{k\alpha\sigma}U_{\alpha}n_{k\alpha\sigma},$ (S24)
@@
VIII Mean-Field Results Away From Filling $\nu=-1$
In the experiment, the $3.65^{\circ}$ tWSe2 system exhibits metallic behavior on both sides as the filling deviating from $\nu=-1$. In this section, we present a preliminary theoretical investigation on the doping dependence of tWSe2 near the SC phase through mean-field analysis, demonstrating that our model can also explain the key experimental features when the filling is away from $\nu=-1$. Theoretically, the commensurate $\sqrt{3}\times\sqrt{3}$ AFM order tendency generally persists over a certain range of filling near $\nu=-1$, leading to a metallic phase since the filling factor (in the enlarged unit cell) deviates from the integer value $\tilde{\nu}=-3$. Meanwhile, the SC tendency also remains robust upon doping away from $\nu=-1$, as the time-reversal symmetry remains intact. The actual ground state is determined by the energetics of these two phases.
Fig. S8(a) illustrates our mean-field results for the filling dependence of SC and AFM orders at $\mathcal{V}_{z}=0$ near $\nu=-1$, indicating that the SC phase persists only within the range $0.9\lesssim|\nu|\lesssim 1.2$, while an AFM metal phase emerges beyond this filling range, qualitatively consistent with the experimental observations. To better understand the filling dependence of these orders, we present the Fermi surface DOS as a function of $\nu$ in Fig. S8(b), where the DOS is significantly enhanced (suppressed) on the less (more) hole-doped side due to its closer (further) proximity to the van Hove singularity. The Fermi surface DOS provides a natural understanding for the strength of SC order, as well as the enhanced AFM order on the less hole-doped side. In contrast, the emergence of AFM order on the more hole-doped side, as shown in Fig. S8(c), arises from the better nesting condition for the $\sqrt{3}\times\sqrt{3}$ AFM order.
Figure S8: (a) The energy gain per hole of the AFM order and the mixed $p_{x}\pm ip_{y}$ and $d_{xy}\mp id_{x^{2}-y^{2}}$-wave SC order. (b) the Fermi surface DOS as a function of filling $\nu$. (c) Fermi surface at $\nu=-1.3$, where the nesting wave vector $\bm{Q}=(0,\pm 4\pi/3a_{M})$ is represented as black arrow.